Find all heronian triangle with one altitude with 40cm
1
Expert's answer
2019-12-25T08:21:48-0500

A Heronian triangle is a triangle that has side lengths and area that are all integers.

Also,


"Area=\frac{1}{2}base\times height\\=\frac{1}{2}base\times 40\\=base\times 20"

For area to be integer the base should be a natural number. Lengths of two other sides of the triangle also should be natural.